Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Munich World Cup 2001

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Munich
Date June 3 to June 10
Disciplines 10m Air Rifle, 50m Prone Rifle

DURC Attendees

Mark Dennehy Head of Delegation

Scores

Alan Lewis 583
Rhona Barry 383 (MQS)

Report

The Munich World Cup was attended this year by 65 nations with nearly 650 shooters competing. Ireland sent a team of two this year, Miss Rhona Barry (who competed in the Atlanta Olympic games) and Alan Lewis (who competed in the Sydney Olympic games). Mark Dennehy, DURC Secretary, was sent to act as Team Manager.

Alan’s competition, mens 50m Prone Rifle, did not go well for him and he missed the cutoff in the elimination round by one place, with a score of 583 (584 being high enough to qualify for the next round).

Rhona did not have an elimination round in her event, women’s 10m Air Rifle. She shot a 383 in qualification, which is a Minimum Qualifying Score (MQS) for the Athens Olympic Games. This makes her the first Irish shooter to qualify for selection to fill a quota place for the games, and also one of the first Irish athletes to do so.

Mark was present in an administrative capacity during the event, and also represented the Irish contingent at the Technical Meeting and the Delegates Meeting during the World Cup.

Misé le meas,

Mark Dennehy, DURC Secretary